OpenStudy (jiteshmeghwal9):

One day John moves 10km to south from his home , then he moves 5km. to his right hand side & again moves 10km to his right & then moves 10km to his left hand side. now how many km should he cover through cycle to reach home ?

OpenStudy (anonymous):

x^2=10^2+5^2

OpenStudy (anonymous):

Solve for x

OpenStudy (jiteshmeghwal9):

x^2=125 x=5sqrt{5}

OpenStudy (anonymous):

Yep

OpenStudy (jiteshmeghwal9):

so this is the answer ????

OpenStudy (jiteshmeghwal9):

@sauravshakya

OpenStudy (jiteshmeghwal9):

i know the answer but this is a reasoning :)

OpenStudy (anonymous):



OpenStudy (anonymous):

got it?

OpenStudy (jiteshmeghwal9):

sorry ! u have made the wrong construction

OpenStudy (anonymous):

U mean the directions

OpenStudy (jiteshmeghwal9):

yup!

OpenStudy (jiteshmeghwal9):

@hartnn @shubhamsrg @experimentX have a try :)

hartnn (hartnn):